# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ 6.2 Causality Analysis
While the descriptive results suggest a concordance between greater senior-level diversity and higher firm performance, correlation alone cannot establish causality. In this section, we formally test whether increases in executive diversity causally impact Tobin’s $Q$, using the methodology developed in Sections 24. Because the “treatment” (crossing a diversity threshold) is not randomly assigned, a naïve estimation of this effect risks bias from selection on unobservables. We therefore implement both a conventional point-estimation approach under strong assumptions and a robust partial-identification approach under minimal assumptions, and compare the findings.
@@
First, we apply an unconditional mean-comparison framework following Angrist and Pischke, (2009). For each candidate diversity threshold $\tau$ (e.g. 5%, 10%, …, 50%, etc.), firms are split into a treated group (above the threshold) and a control group (below the threshold). We then estimate the difference in mean Tobin’s $Q$ between treated and control firms for that threshold. This difference-in-means is a point estimate of the ATE if one assumes mean independence (i.e. that, conditional on crossing the threshold, potential outcomes are the same for treated and control firms on average). We construct simultaneous 95% confidence bands for these ATE estimates across all thresholds in the set M=5,10,15,,90,95M={5,10,15,…,90,95}, applying a Bonferroni or Šidák correction to account for the multiple comparisons. This yields a series of tests for the null hypothesis of no effect at each diversity level, adjusted so that the overall family-wise error rate is 5%. It is important to note that this point-identified approach treats the threshold “treatment” as if random; in practice, firms that surpass a given diversity level could differ systematically from those that do not (for instance, more progressive or better-governed firms might both adopt diverse leadership and perform well for other reasons). As a result, the point estimates of $\delta$ may capture more than the true causal effect of diversity. We use this method as a benchmark, fully aware that its validity hinges on strong assumptions.
@@
We next relax the strong assumptions by employing a partial identification strategy (Manski,, 1990, 2003). Instead of assuming we can precisely identify the counterfactual outcome for each firm, we derive bounds on the possible ATE. Instead of point identification, we partially identify the region in which the average treatment effect $\Re$ lies, as characterized by Eq.(14). We denote this set the identification region $H[\Re_{u}]$ for all $u$ in $\mathcal{M}$, where as noted in Section 2.3, $\mathcal{M}=\{5,10,15,\cdots,90,95\}$ which represents the random diversity thresholds. As previously noted, estimation of Eq. (14) involves latent quantities $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}=1]$ and $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}=0]$, which are not observed but can be bounded by quantities $L^{(k)}$ and $U^{(k)}$. On on hand, we may acknowledge that the extrema of the latent outcomes within the finite sample may not capture the true population extrema (and consequently the true treatment effect interval), in which case we rely on the finite sample hybrid approach. On the other hand, one may argue that since using the full range of outcomes ($\min$ and $\max$) can lead to overly conservative bounds, we also construct Manksi bounds using the (5th, 95th) and (10th, 90th) quantiles of $Y_{ijt}^{(k)}$. Finally, we build a simultaneous joint 95% confidence region for the estimated bounds to make causal inference claims.
@@
Before turning to results, we address some practical implementation details. As noted in Section 3, it is necessary for both the treated and control groups to be non-empty (and sufficiently large) at each threshold to estimate meaningful effects. In our panel, some extreme diversity thresholds (especially very high ones) result in very few firms in one group. We therefore discard threshold levels $\tau$ for which one of the groups contains fewer than 10 observations (approximately, we require at least 10 firm-quarters above and below the threshold). If too many high-$\tau$ values are discarded for a particular subset of the data, that subset is excluded from the threshold analysis due to lack of support. In practice, this means that for some sector-specific analyses we cannot evaluate very high diversity percentages because, for example, no firm in a given sector ever reaches 90% female leadership. Based on this criterion, certain combinations of sector and diversity type are dropped from the causal analysis. In particular, we exclude female leadership in sectors that never approach gender parity (notably the Financials and Energy sectors). These exclusions are a matter of data availability and ensure that the identification regions for ATE do not trivially collapse to a point. All remaining sector clusters and diversity measures satisfy $0<\Pr(Z=1)<1$ at the thresholds of interest, so both treated and control outcomes can be observed in those cases.
@@
We first discuss the overall sample (pooling all industries). The naive point estimates suggest that higher diversity leads to better performance, with a tipping point emerging at a moderate level of female representation. In particular, the Angrist–Pischke mean comparison indicates that once female executives make up roughly one-third of the top team, the estimated ATE on Tobin’s $Q$ becomes positive and statistically significant (the 95% confidence band excludes zero). This is evidenced by a significant jump at about the 30–35% female share in leadership (see Table 6.2, which summarizes the estimated threshold levels at which the treatment effect becomes significant under each method). These point-identification results, however, must be interpreted cautiously, since they assume no unobserved confounding. Indeed, the partial identification analysis reveals that under weaker assumptions the evidence is less definitive. The Manski bounds and our hybrid concATE bands for the overall sample show an increasing trend as diversity rises, but crucially the confidence band for the ATE always includes zero at conventional confidence levels. The estimated bounds exhibit a sigmoid-like shape: at very low diversity levels, the lower bound on the ATE is substantially negative (reflecting the possibility that token diversity could harm performance or that the most homogeneous firms might be unusually strong performers), but this lower bound rises toward zero as diversity increases. We observe an inflection point around approximately 20–25% female representation, beyond which the bounds begin to narrow. This value is close to the “critical mass” threshold theorized by Kanter, (1977) – the point at which a minority group’s representation shifts from tokenism to a more influential presence. Above roughly 25% female share, the worst-case (lower-bound) effect of diversity is no longer hugely negative; it hovers near zero, while the upper-bound effect is positive. Nevertheless, even at the highest levels of diversity observed (e.g. 80–90% female leadership), the 95% joint confidence region for the ATE bounds still straddles zero under the baseline (full-range) Manski scenario. In other words, without stronger assumptions we cannot conclusively assert a positive causal effect in the full sample – the data are consistent with a benefit from diversity, but also with no effect. This underscores the importance of robust inference: what appears significant in the point estimate can become statistically ambiguous once we account for uncertainty about counterfactual outcomes.
@@
Note: This table presents the estimated tipping points—i.e., the random diversity thresholds at which the diversity treatment has a significantly positive effect on Tobin’s $Q$. Cells marked with a (–) indicate cases where significance is not achieved at any of the prescribed thresholds. Rows labeled “N/A” correspond to cases that do not meet the minimum threshold size condition of $\tau_{m}>50$ discussed earlier.
If we incorporate mild additional assumptions by trimming the outcome tails, the partial identification results become slightly more optimistic, though still cautious. For example, imposing that the true outcome lies within the 5th and 95th percentile of observed Tobin’s $Q$ (thereby excluding implausibly extreme counterfactuals) yields somewhat tighter bounds. In this case we find that at high diversity levels (for instance, beyond 60–70% female leadership) the lower bound of the ATE nearly exceeds zero. Using the 10th and 90th percentile restrictions – a stronger assumption that rules out the extreme 10% tails – we even see the lower bound move just above zero for some thresholds. However, these effects are marginal, and at the 95% confidence level the concATE band for the overall sample still does not fully exclude zero for any threshold when using only the weakest (10th/90th) trimming. The overall conclusion is that, in the full sample, the positive relationship between diversity and performance could be causal, but a conservative analysis cannot rule out a zero impact. The naive method’s significant findings (e.g. a female-share tipping point around one-third) may reflect underlying selection bias or favorable unobserved characteristics of diverse firms, since those findings disappear when we allow for more uncertainty.
We also examine the causal effects within certain sector groupings, focusing on cases where sufficient diversity variation exists. Notably, the Cyclicals sector (which includes industries like Consumer Discretionary and Industrials) shows a clearer pattern of a diversity tipping point. In this sector, our hybrid partial-identification band indicates that once women comprise more than about 55% of senior management – i.e. when female leadership surpasses men – the ATE on Tobin’s $Q$ becomes positive and statistically distinguishable from zero. In other words, for cyclically sensitive industries, achieving majority-female leadership is associated with a robust increase in firm value. Interestingly, the naïve point estimate in cyclicals also eventually signals a positive effect of female leadership, but it finds significance at a somewhat lower threshold (around 30% in our data). This discrepancy highlights how the point estimate can give a premature indication of significance by not accounting for potential biases; the concATE band insists on a higher threshold (and thus a larger performance gain) before declaring the effect significant, reflecting a more stringent standard of evidence. Turning to the Growth & Innovation sector (technology and communications firms), female leadership also appears to have a strong effect once it reaches a critical mass. The hybrid confidence band for the female-share ATE in growth industries becomes significantly positive at roughly 50–55% (about equal gender representation). This aligns with the idea that innovative firms may harness diverse perspectives especially well once a balanced team is in place. The point-estimate analysis also reflects a positive effect in this sector, and interestingly it suggests significance starting at approximately the same range (around the 50% threshold), reinforcing the partial-bounds finding in this case.
In more traditional or constrained sectors such as Defensives and Energy, the lack of sufficient high-diversity observations means our method finds no significant effects. In the Defensive industries (e.g. utilities, healthcare), few firms exceeded 40% female leadership during the sample, and accordingly neither the naive nor the robust approach indicates any clear performance gains from diversity — indeed, the concATE bounds always remain wide and centered around zero. In Energy, as noted, diversity levels are generally low and sometimes inversely related to performance in simple correlations. Consistently, we do not find any positive causal effect in Energy firms; if anything, the point estimates for female leadership in Energy were negative (though not significant under bounds).
@@
Figure 6.4: Hybrid and Manski’s Nonparametric Bounds, and Angrist’s Point Estimates - (Significant Results)
Finally, in the Financials sector, female representation never reached high levels so the female diversity effect is not estimable there with our approach (we excluded that case).
In summary, the causality analysis using nonparametric bounds and finite-sample confidence bands paints a more nuanced picture than the raw correlations. The data provide qualified evidence of “tipping points”: in certain high-growth or cyclical environments, reaching a critical mass of diversity (for example, women comprising about half of senior leadership roles) is associated with a reliable increase in firm value. However, under the robust inference approach, these effects emerge at higher thresholds and with less ubiquity than a naïve analysis would suggest. The concATE methodology ensures that any detected effect is robust to heavy-tailed outcomes and potential selection bias, guarding against false positives. Where the conventional method finds significance at lower diversity levels, the bounded analysis often still intersects zero, meaning we cannot rule out an absence of effect without further assumptions.
@@ 7 Concluding Remarks
This paper introduces concATE as a general framework for robust causal inference when point identification is not possible or reliable. By marrying Manski,’s (1990; 2003) nonparametric bounds with finite-sample concentration inequalities, concATE offers researchers a new tool to obtain ATE confidence bands without assuming away heavy-tailed outcomes or requiring strong parametric models. The methodology’s broader relevance lies in its ability to deliver valid inference under minimal assumptions and even with weakly dependent data, thereby guarding against false positives that can arise from conventional point estimates under mis-specified models or overlooked tail risks.
Our empirical findings on workforce diversity illustrate the importance of such rigorous inference. While naïve regressions might suggest that even modest increases in female leadership yield significant gains, the concATE approach paints a more nuanced picture. We find that substantive benefits of gender diversity materialize only once a sufficient representation level is achieved. In practice, this means token diversity (e.g. a lone female or two in senior leadership) is unlikely to drive measurable performance improvement, whereas reaching a critical mass of women in leadership – roughly one-third or more in growth-oriented industries (and higher in others) – is associated with a reliably positive impact on firm value. These conclusions align with the critical mass hypothesis: diversity can boost performance, but only after crossing a threshold that moves an organization beyond tokenism (Kanter,, 1977). By confirming this pattern under stringent inference, our study provides guidance for firms and policymakers – emphasizing that real gains from diversity require either significant numbers of women or alternatively substantial inclusion efforts [^1] We note that an inclusive organizational culture may allow firms to reap performance gains at lower diversity levels than this critical mass. In settings without an inclusive culture, a small number of women leaders often remain marginalized “tokens” with limited influence, but when genuine inclusion is present, even a few women in leadership can contribute meaningfully to performance improvements. It is an area for future research to investigate the relationship between inclusion and diversity. – and also demonstrates how concATE can be applied in other domains to uncover robust causal insights where traditional methods may be misleading.

### v2

@@ 6.2 Causality Analysis
While the descriptive results suggest a concordance between greater senior-level gender diversity and higher firm performance, correlation alone cannot establish causality. In this section, we formally test whether increases in executive diversity causally impact Tobin’s $Q$, using the methodology developed in Sections 24. Because the “treatment” (crossing a diversity threshold) is not randomly assigned, a naïve estimation of this effect risks bias from selection on unobservables. We therefore implement both a conventional point-estimation approach under strong assumptions and a robust partial-identification approach under minimal assumptions, and compare the findings.
@@
First, we apply an unconditional mean-comparison framework following angrist2009mostly. For each candidate diversity threshold $\tau$ (e.g. 5%, 10%, …, 50%, etc.), firms are split into a treated group (above the threshold) and a control group (below the threshold). We then estimate the difference in mean Tobin’s $Q$ between treated and control firms for that threshold. This difference-in-means is a point estimate of the ATE if one assumes mean independence (i.e. that, conditional on crossing the threshold, potential outcomes are the same for treated and control firms on average). We construct simultaneous 95% confidence bands for these ATE estimates across all thresholds in the set $\mathcal{M}={5,10,15,…,90,95}$, applying a Bonferroni or Šidák correction to account for the multiple comparisons. This yields a series of tests for the null hypothesis of no effect at each diversity level, adjusted so that the overall family-wise error rate is 5%. It is important to note that this point-identified approach treats the threshold “treatment” as if random; in practice, firms that surpass a given diversity level could differ systematically from those that do not (for instance, more progressive or better-governed firms might both adopt diverse leadership and perform well for other reasons). As a result, the point estimates of $\delta$ may capture more than the true causal effect of diversity. We use this method as a benchmark, fully aware that its validity hinges on strong assumptions.
@@
We next relax the strong assumptions by employing a partial identification strategy (manski1990nonparametric; manski2003partial). Instead of assuming we can precisely identify the counterfactual outcome for each firm, we derive bounds on the possible ATE. Instead of point identification, we partially identify the region in which the average treatment effect $\Re$ lies, as characterized by Eq.(14). We denote this set the identification region $\mathcal{H}_{\alpha_{u}}[\Re_{u}]$ for all $u$ in $\mathcal{M}$, where as noted in Section 2.3, $\mathcal{M}=\{5,10,15,\cdots,90,95\}$ which represents the random diversity thresholds. As previously noted, estimation of Eq. (14) involves latent quantities $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau_{u})=1]$ and $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau_{u})=0]$, which are not observed but can be bounded by quantities $L^{(k)}$ and $U^{(k)}$. On one hand, we may acknowledge that the extrema of the latent outcomes within the finite sample may not capture the true population extrema (and consequently the true treatment effect interval), in which case we rely on the finite sample hybrid approach. On the other hand, one may argue that since using the full range of outcomes ($\min$ and $\max$) can lead to overly conservative bounds, we also construct Manksi bounds using the (5th, 95th) and (10th, 90th) quantiles of $Y_{ijt}^{(k)}$. Finally, we build a simultaneous joint 95% confidence region for the estimated bounds to make causal inference claims.
@@
Before turning to results, we address some practical implementation details. As noted in Section 3, it is necessary for both the treated and control groups to be non-empty (and sufficiently large) at each threshold to estimate meaningful effects. In our panel, some extreme diversity thresholds (especially very high ones) result in very few firms in one group. We therefore discard threshold levels $\tau$ for which one of the groups contains fewer than 10 observations (approximately, we require at least 10 firm-quarters above and below the threshold). If too many high-$\tau$ values are discarded for a particular subset of the data, that subset is excluded from the threshold analysis due to lack of support. In practice, this means that for some sector-specific analyses we cannot evaluate very high diversity percentages because, for example, no firm in a given sector ever reaches 90% female leadership. Based on this criterion, certain combinations of sector and diversity type are dropped from the causal analysis. In particular, we exclude female leadership in sectors that never approach gender parity (notably the Financials and Energy sectors). These exclusions are a matter of data availability and ensure that the identification regions for ATE do not trivially collapse to a point. All remaining sector clusters and diversity measures satisfy $0<\Pr(Z(\tau_{u})=1)<1$ at the thresholds of interest, so both treated and control outcomes can be observed in those cases.
@@
We first examine the impact of senior-level gender diversity on firm performance in the full sample, comparing a naive point estimate approach to our partial identification methods. Using the unconditional difference-in-means (Angrist and Pischke’s approach), we find that greater female representation is associated with higher Tobin’s $Q$, with an apparent “tipping point” at moderate diversity levels. In particular, once women comprise roughly one-third of the top management team, the naive ATE estimate becomes positive and statistically significant. For example, crossing about 30–35% female leadership is associated with a jump in Tobin’s $Q$ (see Table 6.2, which summarizes the estimated threshold levels at which the treatment effect becomes significant under each method): beyond this threshold the simple treated-control difference excludes zero at the 5% level. This suggests that, under strong assumptions of ignorability, even a moderately gender-diverse leadership team might boost firm market value.
However, when we relax those assumptions, the evidence is less definitive. Manski’s nonparametric bounds, which allow for arbitrary selection and unobserved heterogeneity, remain wide in finite samples, and their associated confidence band always includes zero. At very low diversity levels, the lower bound on the ATE is substantially negative (reflecting the worst-case scenario that “token” diversity could coincidentally occur in poorly performing firms). As the female share increases, this lower bound rises toward zero. We observe an inflection around 20–25% female representation, roughly consistent with kanter1977some’s notion of moving from tokenism to a more influential minority. Beyond that point, the worst-case impact of diversity is no longer severely negative; by around 50% female leadership, the lower bound is near zero and the upper bound is positive. Nevertheless, without additional information about outcome limits, even at the highest diversity levels observed (e.g. 80–90% female), the 95% confidence region for the ATE still straddles zero. In other words, under minimal assumptions the data do not allow us to conclusively rule out no effect (or even a small negative effect) for the overall sample. This highlights how misleading the precise naive estimate can be: what appears as a clearly positive effect with a simple mean comparison becomes statistically ambiguous once we account for uncertainty about counterfactual outcomes.
@@
Note: This table presents the estimated tipping points—i.e., the random diversity thresholds at which the diversity treatment has a significantly positive effect on Tobin’s $Q$. Cells marked with a (-) indicate cases where significance is not achieved at any of the prescribed thresholds. Rows labeled “N/A” correspond to cases that do not meet the minimum threshold size condition of $\tau_{m}>50$ discussed earlier.
Imposing mild outcome bounds yields somewhat tighter inference. If we assume, for instance, that Tobin’s $Q$ outcomes are effectively constrained within the central 90–95% of the observed sample range (excluding extreme tail realizations), the identified ATE interval narrows. Under these plausible restrictions, the partial identification bounds move inward: the lower bound is higher (less negative) and the upper bound lower (less positive) than the unbounded Manski case. As a result, the concATE confidence band becomes more optimistic at high diversity levels. For example, using the 5th and 95th percentiles of $Q$ as rough bounds, we find that at very high diversity (above about 60% women in leadership) the lower bound on the ATE is nearly zero or slightly positive. With an even tighter 10th–90th percentile restriction, the lower bound actually rises above zero at some thresholds. These results hint that a real positive effect may emerge once diversity is sufficiently high: with female leadership above roughly two-thirds, even the worst-case impact is likely to be zero. However, we emphasize that even under these trimmed-outcome assumptions, the joint 95% confidence band for the ATE barely excludes zero. In the full sample, no diversity threshold produces a completely robust positive effect at the 5% significance level unless one accepts some outcome-range assumptions. Thus, our most cautious conclusion for the overall dataset is that greater gender diversity could improve firm value, but the evidence is not statistically conclusive under minimal assumptions. The contrast between the naive point estimate (significant at $\sim$30% diversity) and the conservative bounds (no significance without assumptions) underscores the importance of conservative, rigorous inference: smaller apparent gains may reflect unobserved biases or heavy-tailed outcomes rather than true causal effects.
Sector-Specific “Tipping Points”
We next investigate whether the diversity–performance relationship exhibits stronger effects in particular types of firms. To explore this, we apply an identical analysis within more homogeneous sector groupings. In each case we report the threshold at which the concATE band indicates a significant effect, and compare it to the naive and classical bounds results. This reveals several interesting findings.
First, in high-growth, innovation-intensive industries, we find clear evidence of a diversity tipping point. Firms in these sectors show relatively high variance in leadership composition, with some approaching gender-balanced teams. The unconditional mean comparison suggests a positive effect of diversity that becomes sizeable at upper diversity levels. However, due to the smaller sample of firms in this category, the naive threshold for significance is somewhat high: only at nearly half female representation does the simple difference in Tobin’s $Q$ become significant. Our robust analysis confirms and sharpens this finding. The concATE confidence band for the average treatment effect in Growth & Innovation firms first excludes zero at approximately 55% female representation in senior roles. In other words, once a firm’s top team is roughly half women, we can confidently assert a positive causal impact on market valuation in this sector. Below that threshold, the partial-identification interval still includes zero, meaning the effect cannot be distinguished from zero (i.e. while we cannot assert a positive effect, we can rule out a negative effect) with high confidence. Notably, the estimated ATE grows larger as diversity increases beyond 55%; for firms that actually achieve gender-balanced or women-majority leadership, even the lower bound of the effect is distinctly above zero. This pattern aligns with the idea that innovative companies reap substantial benefits from diverse perspectives only after achieving a critical mass of diversity. Before that point, female voices may be too diluted to change organizational outcomes, but around parity their influence on decision-making and the innovation climate becomes strong. It is encouraging that both the naive method and the more rigorous concATE method point to a similar threshold in these sectors (around 50–55% female): this convergence suggests the result is not merely an artefact of assumptions. In sum, for Growth & Innovation firms we find a statistically significant positive causal effect of diversity emerging at just over half women in leadership.
Second, for Defensive sectors (Healthcare/Staples/Utilities), traditionally “stable” industries that historically have lower female leadership representation, fewer firms in our sample reach high diversity levels. A naive analysis indicates that even moderate diversity might help performance: the difference-in-means suggests an uptick in Tobin’s $Q$ once the female share surpasses roughly 40% in these sectors. Indeed, raw correlations in the Defensive group are positive, hinting that more diverse leadership teams tend to coincide with slightly higher $Q$ ratios. However, our robust inference reveals that the bar for significance is higher in this context. The concATE confidence band does not exclude zero until female representation reaches around 60% or more in Defensive-sector firms. In other words, only when women form a substantial majority of top management do we find a clear positive effect on firm value with 95% confidence. This implies that at the 60% threshold there are true performance gains from diversity. One interpretation is that these traditional industries require a larger critical mass to overcome legacy cultures and realize the advantages of inclusion. When women remain a small minority (say 20–30%), they may lack the influence or psychological safety needed to affect corporate strategy, yielding no measurable gain. By contrast, if a firm reaches 60% female leadership (an uncommon achievement), it likely reflects deep organizational changes that unlock diversity’s benefits (e.g. improved problem-solving, stakeholder alignment, or innovation even in mature markets). Thus, for Defensive sectors our findings suggest a delayed tipping point: meaningful performance improvements emerge only at a high level of representation, higher than in fast-paced growth industries. This result underscores how the required “critical mass” can vary by context.
The remaining sectors (Cyclical, Financial, Energy) either showed no robust threshold within our data range or could not be rigorously analysed due to limited support. Cyclical sectors (e.g. Consumer Discretionary, Industrials) have intermediate diversity levels. The naive analysis in cyclicals suggested a possible positive effect emerging at about 30% female leadership (similar to the overall sample). Yet, using our more cautious approach, we found that the confidence intervals for the ATE in cyclicals still included zero at all feasible thresholds. In short, we cannot confirm a statistically significant benefit even if the point estimates are positive. We can, however, rule out a negative impact. This does not mean diversity has no effect in cyclical firms, but rather that the data do not provide high-confidence evidence of an effect under minimal assumptions. It is possible that unobserved factors or heavy-tailed outcomes obscure the impact in this mixed group of industries.
For Financial firms, we were unable to identify a tipping point because virtually none of the sampled banks or insurers exceeded 45–50% female leadership during the study period. Since testing a threshold requires some treated and control firms on either side, the lack of any instances of very high diversity meant we could not apply our sequential threshold test in Financials. Interestingly, the correlation between diversity and Tobin’s $Q$ in finance was negative in earlier years and then became positive toward the end of our sample (as noted in our descriptive analysis), suggesting a shifting dynamic. Our method would need a longer horizon or more variation to pin down where a critical mass effect might occur in finance, if at all.
The Energy sector remains an outlier. Energy companies not only had the lowest levels of female leadership (maxing out around 50% in our data, with most far lower), but they also exhibited a negative raw correlation between diversity and performance. This negative association could reflect reverse causality or industry-specific factors. For instance, struggling energy firms might appoint more women to leadership in response to external pressures, creating a spurious negative link. In any case, our partial identification analysis did not find any significant positive effect of diversity in Energy. Even at the highest observed female share (just above 50%), the ATE bounds encompass zero and even negative values. Thus, we find no evidence of a beneficial tipping point in Energy firms. We caution that this does not prove diversity harms performance in energy—only that, given the data and minimal assumptions, we cannot confirm any uplift. It is a reminder that the advantages of diversity may not be universal and could depend on complementary organizational changes, such as a culture of inclusion.
@@
Figure 6.4: Hybrid and Manski Nonparametric Bounds
Note: Shaded regions represent the nonparametric upper and lower bounds on the average treatment effect. Solid lines denote the midpoint between the upper and lower bounds.
In summary, the causality analysis using nonparametric bounds and finite-sample confidence bands paints a more nuanced picture than the raw correlations. The data provide qualified evidence of “tipping points”: in certain high-growth or defence sectors, reaching a critical mass of diversity (for example, women comprising about half of senior leadership roles) is associated with a reliable increase in firm value. For the remaining sectors (with the exception of the energy sector), we can conclude that diversity does not change firm value (i.e. it does not have any negative effects).
@@ 7 Concluding Remarks
This paper introduces concATE as a general framework for robust causal inference when point identification is not possible or reliable. By marrying Manski’s nonparametric bounds with finite-sample concentration inequalities, concATE offers researchers a new tool to obtain ATE confidence bands without assuming away heavy-tailed outcomes or requiring strong parametric models. The methodology’s broader relevance lies in its ability to deliver valid inference under minimal assumptions (even with weakly dependent data), thereby guarding against false positives that can arise from conventional point estimates under misspecified models or overlooked tail risks.
Our empirical findings on workforce diversity illustrate the importance of such rigorous inference. While naive regressions might suggest that even modest increases in female leadership yield significant gains, the concATE approach paints a more nuanced picture. We find that substantive benefits of gender diversity materialize only once a sufficient representation level is achieved. In practice, this means token diversity—for example, a lone woman or two in senior leadership—is unlikely to drive measurable performance improvement. By contrast, reaching a critical mass of women in leadership (roughly half or more in growth-oriented industries, and a clear majority in others) is associated with a reliably positive impact on firm value. These conclusions align with the critical mass hypothesis: diversity can boost performance, but only after crossing a threshold that moves an organization beyond tokenism (kanter1977some).
By confirming this pattern under stringent inference, our study provides guidance for firms and policymakers. It emphasizes that real gains from diversity require either significant numbers of women or, alternatively, substantial inclusion efforts. Indeed, an inclusive organizational culture may allow firms to reap performance gains at lower diversity levels than this critical mass. In settings without an inclusive culture, a small number of women leaders may remain marginalized “tokens” with limited influence. However, when genuine inclusion is present—for example, leadership practices that actively include and value minority voices—even a few women in leadership can contribute meaningfully to performance improvements (nishii2013benefits; roberson2006disentangling). This perspective is consistent with evidence that diversity alone is not sufficient and must be accompanied by inclusion to realize its full benefits (almeida2024diversity; josten2025makes). Investigating the relationship between inclusion and diversity outcomes remains an important area for future research. Finally, our application also demonstrates how concATE can be applied in other domains to uncover robust causal insights where traditional methods may be misleading.